OpenStudy (anonymous):

how do you solve for t? 1*3.34*10^5 +(T-0)(4160) = 4160 (50-T)(9)

OpenStudy (anonymous):

is the answer 0?

OpenStudy (phi):

is the problem \[ 1 \cdot 3.34 \cdot 10^5 +(T-0)(4160) = 4160\cdot 9\cdot (50-T) \] ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

\[ 334000 +4160T = 1872000 - 37440 T \\ 41600 T = 1538000 \\ T= 36.971 \]

OpenStudy (anonymous):

ohh thanks

OpenStudy (anonymous):

do you happen to know anything about physics because now im questioning the use of that equation for the problem?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

1 kg of ice at 0° C is mixed with 9 kg of water at 50° C (The latent heat of ice is 3.34x105 J/kg and the specific heat capacity of water is 4160 J/kg). What is the resulting temperature?